OpenStudy (akohl):

@Abhisar A red car of mass m is heading north (direction 0°). It collides at an intersection with a yellow car of mass 1.3m heading east (direction 90°). Immediately after the collision, the cars lock together and travel at in direction 42°. What is the speed of the yellow car just before the collision? 24 m/s 28 m/s 11 m/s 17 m/s

OpenStudy (akohl):

The velocity after collision is 20m/s

OpenStudy (abhisar):

Initial velocity of red car is not given ?

OpenStudy (akohl):

No

OpenStudy (akohl):

The red car mass is variable m and yellow is 1.3*m

OpenStudy (abhisar):

@Astrophysics

OpenStudy (astrophysics):

Mass 1.3m lol mhm o_O, well do you have any idea on how to start it? Anything that comes in mind when it says collision?

OpenStudy (akohl):

Mass *v is the momentum

OpenStudy (astrophysics):

Yes this question deals with momentum, so first thing, draw a picture out :), and what for 1.3m mass mean? Oh I see the yellow is 1.3* the mass of red car hehe. The wording! Anyways any attempt at the drawing?

OpenStudy (akohl):



OpenStudy (anonymous):

Seems like there is some information missing when it says " the cars lock together and travel AT ... IN direction 42 degrees.

OpenStudy (akohl):

20m/S

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @akohl The velocity after collision is 20m/s \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

So i think we will have to apply law of conservation of momentum ?

OpenStudy (akohl):

YEA mass times velocity

OpenStudy (astrophysics):

\[\sum p _{i}= \sum p _{f}\] does this ring a bell? (remember these are vector quantities)

OpenStudy (abhisar):

But initial velocity of red ball is not given.

OpenStudy (astrophysics):

Newtons third law basically :P

OpenStudy (astrophysics):

Oh I see

OpenStudy (akohl):

cant add vectors without magnitude

OpenStudy (astrophysics):

Hey, do you have an image of the full question?

OpenStudy (abhisar):

That's wr i got struck !

OpenStudy (akohl):

no there wasnt one

OpenStudy (astrophysics):

Or can you just post wherever you got your question from, I have doubts haha.

OpenStudy (akohl):

I cant its off a worksheet from my summer school teacher lol

OpenStudy (abhisar):

*stuck

OpenStudy (astrophysics):

If we had an initial velocity, this would be cake :p

OpenStudy (anonymous):

 u will get \[V _{1}=2.3 V\] \[V _{2}= \frac{ 2.3 }{ 1.3 }V\] and then use the conservation of kinetic energy and substitute with v1 and v2 to get V \[\frac{ 1 }{ 2 }mV _{1}^{2}+\frac{ 1.3 }{ 2 }mV _{2}^{2}=\frac{ 2.3 }{ 2 }mV ^{2}\]

OpenStudy (anonymous):

oh noo i forgot to write the values cos 45 write it and continue.

OpenStudy (anonymous):

momentum is vector while energy isn't.